5(b) A bond with a face value of $500,000 pays quarterly interest of 1.5% per each period. be willing to pay for this bond today if the next interest payment is due now and you want earn 8% compounded quarterly on your money? Twenty interest payments remain before the bond matures. How much would you
Answer #1

Answer:

Face value P=$500000

Coupon rate =1.5%

Coupon C=1.5%*500000=$7500

r=8%/4=2%

Since interest payment is due now

Price of Bond=C+C*(1-(1+r)^-19)/r+P/(1+r)^19

Price of Bond=7500+7500*(1-(1+2%)^-19)/2% + 500000/(1+2%)^19=$468303.84

> ANSWER IS NOT RIGHT, FIX IT

Szoks Balacs Mon, Feb 14, 2022 6:43 PM